OpenStudy (anonymous):

Find the hydrogen concentration 2,4 = -log[H]

sam (.sam.):

\[2.4 = -\log[H]\] \[-2.4= \log_{10}[H]\] Using logarithm properties, \(log_ab=c \Leftarrow \Rightarrow b=a^c \) \[[H]=10^{-2.4}\] Solve